### v1

# Context-Driven Performance Modeling for Causal Inference Operators on Neural Processing Units

Abstract
The proliferation of large language models (LLMs) has driven demand for long-context inference on resource-constrained edge devices. However, deploying these models on Neural Processing Units (NPUs) presents significant challenges due to the architectural mismatch: quadratic complexity of standard attention mechanisms conflicts with memory and compute patterns of edge accelerators. This paper presents a comprehensive performance analysis of various causal inference operators on a modern NPU. We benchmark standard quadratic attention against several sub-quadratic alternatives, including structured state-space and linear attention models. Our analysis reveals that while sub-quadratic methods offer superior scalability, they introduce distinct computational bottlenecks on the NPU’s specialized execution units. We identify that quadratic attention becomes severely memory-bound, suffering from cache inefficiency and pipeline stalls exceeding 95% at long contexts. In contrast, sub-quadratic models can become compute-bound on programmable vector cores. These findings provide critical insights for the co-design of hardware-aware models and optimization strategies to enable on-device AI inference with long-contexts.
I Introduction
Long-context inference has become essential for document understanding, conversational AI, and real-time decision systems [1, 2]. Applications from medical record analysis to legal contract review increasingly demand processing of 100K+ token sequences. While cloud solutions can handle these workloads, privacy concerns, latency requirements, and operational costs are driving deployment to edge devices. Modern edge platforms now integrate Neural Processing Units (NPUs)—specialized accelerators that deliver exceptional efficiency for data-parallel operations but face severe constraints: limited scratchpad memory (typically 2-4MB), systolic execution patterns, and strict power budgets.
Transformer-based models like Llama deliver state-of-the-art quality but require quadratic computation and linear memory growth with context length [3]. At just 16K tokens, the Key-Value cache consumes over 768MB—more than 30x the capacity of leading NPUs. State-space models (SSMs) like Mamba offer linear scaling but underutilize NPU parallelism due to their recurrent nature during autoregressive decoding inference [4]. This fragmentation forces practitioners into hardware-specific compromises, limiting adoption of long-context capabilities where they’re needed most: on-device.
Recent works have looked into hybrid architectures or ways towards creating a general class of operators for causal inference [5, 6, 7]. While these efforts advance the theoretical understanding, a critical gap remains in characterizing how these diverse operators perform on real-world, resource-constrained hardware like NPUs. The architectural trade-offs made by these models—such as prioritizing data locality, computational regularity, or state compression—have profound, yet unquantified, implications for on-device performance. This paper bridges that gap by providing a comprehensive performance analysis and modeling of causal inference operators on a modern NPU. We move beyond theoretical complexity and provide a rigorous empirical study of latency, throughput, and hardware utilization, culminating in a Roofline analysis that explains the performance landscape.
I-A Our Contributions
This work provides a detailed characterization of attention mechanisms on edge NPUs. Our key contributions are:
  • A comprehensive performance benchmark of quadratic and sub-quadratic attention operators on a real-world NPU, analyzing latency, throughput, and pipeline efficiency.
  • Identification and analysis of critical performance bottlenecks, demonstrating that quadratic attention becomes memory-bound due to cache inefficiency while sub-quadratic models become compute-bound on specialized vector units.
  • A quantitative Roofline performance model that explains the performance limitations of each operator class in terms of the NPU’s architectural constraints.
  • Actionable insights for co-designing hardware-aware models and compiler optimizations to enable efficient long-context inference on the edge.
II Background
II-A Fundamental Tradeoff for Long-Context Inference
Autoregressive sequence modeling faces dual constraints: memory complexity for context retention and computational complexity for state evolution. For sequence length $N$ and model dimension $D$, we observe:
$\underbrace{\text{Memory}}_{\text{Context}}\sim O(N\cdot D),\quad\underbrace{\text{Compute}}_{\text{Inference}}\sim O(N^{2}\cdot D)$ (1)
Prefill Phase: Computes initial context representation:
$\mathcal{C}=f_{\theta}(X_{1:N})\quad\begin{cases}\text{Attention: }\mathcal{C}=\{K,V\}_{1:N}\\ \text{SSM: }\mathcal{C}=h_{N}\end{cases}$ (2)
Decode Phase: Updates state incrementally:
$y_{t},\mathcal{C}_{t}=g_{\theta}(x_{t},\mathcal{C}_{t-1})$ (3)
Memory-State Tradeoff. Architectures balance expressiveness against hardware constraints as we show in Figure 1:
  • Attention-based models (Llama): Maintain explicit KV cache ($O(N\cdot D)$ memory) enabling rich context access.
  • State-space models (Mamba): Compress context to fixed-size state $h_{t}$ ($O(D)$ memory) at additional computational cost.
We introduce this tradeoff not only to compare model classes, but also because it forms the architectural basis for the structured operators analyzed in this paper. Many recent causal inference variants—such as Toeplitz, Retentive, and Fourier—draw inspiration from state-space models, embedding recurrent or convolutional priors directly into attention layers [8, 4, 5, 9, 10]. These hybrid mechanisms blend the long-range capacity of traditional attention with the computational efficiency and inductive structure of SSMs, making the memory-state tradeoff a key element for evaluating modern causal operators on resource-constrained hardware platforms such as NPU.
Figure 1: Differences in persistent memory and layer-wise dataflow for Attention-based Llama vs. SSM-based Mamba.
II-B Heterogeneous Edge Platform with Neural Processing Units (NPU)
State-of-the-art heterogeneous edge platforms—such as Intel Core Ultra processors (Intel AI PC)—combine traditional multi-core CPUs with specialized accelerators, including Graphics Processing Units (GPUs) and Neural Processing Units (NPUs). These diverse compute elements are integrated into a single System-on-Chip (SoC) architecture, featuring a unified system memory that facilitates seamless communication and efficient data sharing across all cores.
Figure 2: NPU dataflow architecture with processing elements (PEs) and accumulator hierarchy. Note the absence of high-bandwidth memory for persistent state storage.
Each of the three compute cores is designed to exploit specific types of workloads, leveraging their distinct system architectures and local memory hierarchies. The classic multi-core CPU, equipped with a three-level cache hierarchy, excels at handling general-purpose, logic-intensive workloads. GPUs feature Xe cores, each comprising multiple vector engines that are optimized for high data parallelism. Every Xe core includes a Shared Local Memory (SLM) block, accessible by all its vector engines to facilitate efficient intra-core communication. In NPUs, compute acceleration is driven by the Data Path Unit (DPU), comprising of a PE array which employs a structured systolic array architecture to perform operations such as matrix multiplication with minimal data movement. NPUs also incorporate Streaming Hybrid Architecture Vector Engines (SHAVE), which support parallel execution of general-purpose tasks and activation function engine to efficiently compute activations. To further enhance efficiency, Direct Memory Access (DMA) engines are integrated into the NPU, enabling high-throughput data transfers from the global shared system memory to the local, software-managed cache. Due to their controlled data movement and predictable, statically scheduled execution flow, NPUs achieve higher energy efficiency for AI workloads compared to GPUs, which involve complex, dynamic execution patterns and data transfers. This energy efficiency is a critical requirement in resource-constrained edge computing environments, where power limitation is a significant design consideration.
II-C Causal Operators for Attention
Figure 3: From left to right, top to bottom are Full Causal, Toeplitz, Fourier, Retentive Decay, Semiseparable, and Linear Structured attention masks.
We describe in further detail which Causal Attention Masks we select to comprehensively survey and perform performance model analysis of a more general class of causal matrices seen in Figure 3.
Full Causal Mask: The standard causal attention mechanism prevents tokens from attending to future positions using a triangular mask:
$\text{Attention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}+M\right)V$
where $M$ is defined as:
$M_{ij}=\begin{cases}0&i\geq j\\ -\infty&\text{otherwise}\end{cases}$
This ensures position $i$ only attends to positions $j\leq i$.
Linear Attention: Replaces the softmax with kernel functions to enable linear complexity:
$\text{LinearAttention}(Q,K,V)=\phi(Q)(\phi(K))^{T}V$
where $\phi$ is a kernel function (in our implementation, low-rank projections). This reduces complexity from $O(N^{2})$ to $O(N)$.
Toeplitz Structured Attention [11]: Constrains the attention matrix to have constant diagonals:
$W_{ij}=\gamma^{|i-j|}$
$\text{ToeplitzAttention}=\text{softmax}(QK^{T}\odot W)V$
where $\gamma$ is a decay factor. This models position-based decay patterns.
Fourier Structured Attention [12]: Leverages the convolution theorem to compute attention in frequency domain:
$Q_{\omega}$ $=\mathcal{F}(Q)$
$K_{\omega}$ $=\mathcal{F}(K)$
$V_{\omega}$ $=\mathcal{F}(V)$
FourierAttention $=\mathcal{F}^{-1}(Q_{\omega}\odot\overline{K_{\omega}}\odot V_{\omega})$
where $\mathcal{F}$ and $\mathcal{F}^{-1}$ are Discrete Fourier Transform (DFT) and Inverse DFT (IDFT) operations, respectively. This enables a computation complexity of $O(N\log N)$.
Retentive Attention [8]: Introduces a decay mechanism that assigns exponentially decreasing weights to tokens based on relative position:
$W_{ij}=\begin{cases}\gamma^{i-j}&i\geq j\\ 0&\text{otherwise}\end{cases}$
$\text{RetentiveAttention}(Q,K,V)=\text{softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\odot W\right)V$
where $\gamma\in(0,1)$ is the decay rate. This maintains causality while modeling recency bias in attention weights. Retentive attention provides tunable context compression and exhibits hardware-friendly diagonal structure.
III Microbenchmarking Kernels
III-A Experimental Setup
We conduct a series of microbenchmarks to evaluate the performance of various attention mechanisms on a Neural Processing Unit (NPU). The objective is to analyze device utilization and latency as a function of context length. All experiments are executed on a system with the hardware specifications detailed in Table I. The NPU integrates a Digital Signal Processor (DSP) for control flow, a Data Path Unit (DPU) with a systolic array for matrix multiplication, programmable SIMD vector processors (SHAVE cores), and a high-bandwidth Direct Memory Access (DMA) engine.
TABLE I: Hardware Specifications
Component Specification Relevance
CPU 16 cores (8P + 8E) Control Logic
NPU 10 TOPS @ 35W Systolic Array Acceleration
DPU (PE Array) $\times128$ INT8 Matrix Multiplication
Scratchpad 4 MB Persistent State Storage
DMA Bandwidth 64 GB/s Data Movement
SHAVE Cores 8 @ 1.4 GHz Element-Wise Operations
Memory 32 GB LPDDR5X Global Buffer
III-B Device Utilization Analysis
To understand the performance bottlenecks, we profile the execution time spent on the NPU’s primary components: the DPU, DMA, and SHAVE cores. Table II presents the utilization breakdown for Fourier State-Space Attention (FSA) and Decayed Recurrent Attention (DRA), which serves as a proxy for retentive decay mechanisms.
For FSA, the workload is initially DPU-bound at shorter context lengths. However, as the context grows beyond 512 tokens, data movement becomes the dominant factor, and the model becomes DMA-bound. This is primarily due to the $\operatorname{concat}$ operations required to manage the state, which saturate the DMA engine’s bandwidth.
In contrast, DRA exhibits a different bottleneck. While initially compute-bound on the DPU, at context lengths of 1024 and greater, the workload transitions to being SHAVE-bound. The $\operatorname{softmax}$ and element-wise multiply operations, which are offloaded to the programmable SHAVE cores, become the most time-consuming parts of the computation at long contexts.
TABLE II: Device Utilization Breakdown (%). At long contexts, FSA becomes DMA-bound while DRA becomes SHAVE-bound.
Model Context DPU (%) DMA (%) SHAVE (%) Bottleneck
Fourier 128 56.4 23.1 20.5 DPU
256 60.8 25.3 13.9 DPU
512 47.2 46.9 5.9 DMA / DPU
1024 46.6 48.9 4.5 DMA
2048 46.2 52.5 1.2 DMA
4096 48.4 51.3 0.3 DMA
8192 61.1 38.9 0.0 DPU
Retentive 128 68.4 0.0 31.6 DPU
256 64.9 0.0 35.1 DPU
512 61.9 0.0 38.1 DPU
1024 34.9 0.0 65.1 SHAVE
2048 24.6 0.0 75.4 SHAVE
4096 28.1 0.0 71.9 SHAVE
8192 23.6 0.0 76.4 SHAVE
Figure 4 illustrates how NPU utilization shifts with context length. Fourier transitions from DPU- to DMA-bound, while Retentive becomes increasingly SHAVE-bound due to element-wise $\operatorname{softmax}$ overhead.
Figure 4: Utilization of NPU subcomponents (DPU, DMA, SHAVE) for Fourier and Retentive attention mechanisms as a function of context length (log scale). Fourier transitions from DPU-bound to DMA-bound at longer contexts, while Retentive attention becomes increasingly SHAVE-bound due to $\operatorname{softmax}$ and element-wise operations.
III-C Scaling of Latency with Context Length
We measure the end-to-end latency of four attention mechanisms—Fourier Structured Attention (FSA), DRA, Toeplitz Structured Attention (TSA), and Causal Linear Attention (CLA)—across a range of context lengths from 128 to 8192. The results, summarized in Table III, reveal distinct scaling properties for each model.
TABLE III: Latency scaling (ms) as a function of context length for four attention operators.
Context Length Fourier Retentive Toeplitz Linear
128 0.39 0.19 0.06 0.09
256 0.79 0.37 0.08 0.13
512 2.54 0.97 0.11 0.24
1024 6.50 2.52 0.18 0.44
2048 15.24 10.04 0.35 0.72
4096 45.69 39.52 0.59 1.52
8192 347.79 85.41 1.01 3.16
TSA and CLA demonstrate highly efficient, sub-quadratic scaling, with only a marginal increase in latency even at very long contexts. DRA exhibits a more pronounced, near-linear growth in latency, which aligns with its shift to being compute-bound on the SHAVE cores. FSA shows the most significant latency increase, scaling poorly at large context lengths due to its reliance on both DPU-intensive computations and DMA-bound data movement, confirming it as the least scalable of the methods that were benchmarked.
Figure 5: Latency scaling of four causal inference operators with respect to context length (log scale). Toeplitz and Linear exhibit near-linear latency growth, while Fourier scales poorly at long contexts due to DPU and DMA bottlenecks.
III-D Performance Scaling and Bottleneck Analysis
We analyze the performance of each kernel by examining latency, throughput, pipeline efficiency, and memory access patterns across varying context lengths. The results, summarized in Tables IV and V, reveal significant architectural bottlenecks for quadratic-time algorithms when deployed on the NPU.
As shown in Table IV, the latency of standard Causal Attention Masking scales quadratically with the sequence length ($N$). This leads to a dramatic drop in throughput, with Causal processing only 4 operations per second at a context of 8192. In contrast, sub-quadratic methods like TSA and Linear Attention maintain low latency and high throughput, demonstrating their suitability for long-context applications on this hardware.
TABLE IV: Latency and throughput scaling at short ($N=512$) and long ($N=8192$) contexts.
Operator Latency (ms) Throughput (ops/s)
$N=512$ $N=8192$ $N=512$ $N=8192$
Causal 4.21 251.41 237 4
Retentive 3.10 45.10 322 22
Fourier 1.59 170.50 631 6
Linear 0.30 3.81 3333 263
Toeplitz 0.75 5.10 1330 196
The underlying cause of this performance degradation is revealed in Table V. At a context length of 8192, Causal and Retentive attention mechanisms suffer from extremely high pipeline stall rates (96.7% and 94.8%, respectively). This indicates that the NPU pipeline is mostly idle, waiting for data to be fetched from memory. The pull stage, responsible for data retrieval, is the primary contributor to these stalls. This is further corroborated by the poor cache efficiency scores for these models (7.7% for Causal and 28.1% for Retentive), which signify a low degree of data reuse and frequent, costly access to main memory. In contrast, the more structured access patterns of Linear and Toeplitz attention lead to much higher cache efficiency and significantly fewer pipeline stalls.
TABLE V: Efficiency metrics at long context lengths. Stall and cache values are percentages; reuse is in milliseconds.
Operator Context ($N$) Stall (%) Cache Efficiency (%) Reuse (ms)
Causal 8192 96.7 7.7 119.92
Retentive 8192 94.8 28.1 25.62
Fourier 4096 95.2 28.6 24.94
Linear 8192 55.2 83.8 1.94
Toeplitz 4096 36.4 87.9 1.38
Figure 6: Efficiency metrics across causal operators at long context lengths. Stall and cache efficiency are shown as bars (left axis), while state reuse latency is plotted as a line (right axis). Lower stall and reuse values with higher cache efficiency indicate better hardware utilization.
III-E Impact of State Dimension
To assess the impact of model parameters on performance, we benchmarked several kernels at a fixed context length ($N=4096$) while increasing the state dimension ($d_{\text{state}}$) from the default of 16 to 128. As detailed in Table VI, increasing the state dimension leads to a predictable rise in latency across all tested models. Notably, Fourier is the most sensitive, with its latency increasing by over $10\times$, highlighting its high computational cost with respect to the state size. Linear and Toeplitz also show increased latency but remain significantly more efficient, confirming that their performance is a function of both sequence length and model dimension.
TABLE VI: Latency impact of increasing state dimension ($d_{\text{state}}$) at fixed context length $N=4096$.
Operator $d_{\text{state}}=16$ (ms) $d_{\text{state}}=128$ (ms)
Linear 2.39 3.37
Toeplitz 0.65 2.73
Fourier 15.50 56.82
IV Performance Modeling and Roofline Analysis
To quantify the fundamental hardware limitations of causal operators on NPUs, we develop a roofline performance model that incorporates effective hardware ceilings based on our performed measurements on the NPU hardware, across several experiments. While theoretical peaks provide an upper bound (10 TOPS compute, 64 GB/s bandwidth), our characterization reveals that architectural overheads limit achievable performance to just 5% of nominal values. This critical insight forms the basis of our analysis.
IV-A Effective Hardware Ceilings
Through microbenchmarking, we establish realistic performance bounds:
  • Effective Compute Ceiling ($\pi_{\text{eff}}$): 500 GOP/s (5% of 10,000 GOP/s theoretical)
  • Effective Bandwidth Ceiling ($\beta_{\text{eff}}$): 3.2 GB/s (5% of 64 GB/s theoretical)
  • Compute-Memory Inflection: $I_{\text{crit}}=\pi_{\text{eff}}/\beta_{\text{eff}}\approx 156$ Ops/Byte
IV-B Operator Intensity Characterization
For each causal operator at $N=4096$, $d_{h}=64$ (16-bit precision):
TABLE VII: Operational intensity and measured performance at $N=4096$, $d_{h}=64$ (16-bit precision).
Operator Intensity (Ops/Byte) Measured (GOP/s) Theoretical Bound (GOP/s)
Full Causal 61.13 21.4 $3.2\times 61.13=195.6$
Retentive 50.00 53.5 $3.2\times 50=160$
Toeplitz 25.00 12.2 $3.2\times 25=80$
Linear 16.00 14.0 $3.2\times 16=51.2$
Fourier 15.00 0.34 $3.2\times 15=48$
IV-C Roofline Analysis
Figure 7 plots measure performance against operational intensity, revealing severe hardware under utilization:
$10^{1}$$10^{1.3}$$10^{1.7}$$10^{2}$$^{-1}1001$$^{0}$$10^{1}$$10^{2}$$10^{2.7}$$10^{2.19}$Causal (4.3%)Linear (27.3%)Fourier (0.7%)Retentive (33.4%)Toeplitz (15.2%)Operational Intensity II [Ops/Byte]Performance $P$ [GOP/s]Roofline Model with Effective CeilingsEffective Memory CeilingEffective Compute Ceiling
Figure 7: Roofline model revealing severe hardware under utilization. Quadratic operators achieve <<5% of their compute ceiling despite high intensity, while linear methods approach 27% of their memory-bound limit.
IV-D Key Insights
The roofline analysis reveals three fundamental limitations:
[…]
  • Sub-Quadratic Operators Are Bandwidth Limited: Linear attention achieves 27% of its memory-bound limit (14/51.2 GOP/s), constrained by DMA bandwidth. Fourier attention performs worst (0.34/48 GOP/s, 0.7%) due to FFT overheads that violate NPU execution assumptions.
[…]
IV-E Efficiency Analysis
TABLE VIII: Hardware utilization metrics at $N=4096$.
Operator Pipeline Stall (%) Cache Efficiency (%) Compute Utilization (%)
Full Causal 96.7 7.7 4.3
Retentive 94.8 28.1 33.4
Toeplitz 36.4 87.9 15.2
Linear 55.2 83.8 27.3
Fourier 95.2 28.6 0.7
Figure 8: Breakdown of hardware utilization at $N=4096$. Full Causal and Fourier operators exhibit high pipeline stall rates with minimal compute utilization. In contrast, Toeplitz and Linear demonstrate better cache efficiency and improved utilization, reflecting tighter memory-compute coupling.
This model proves that memory access patterns, not theoretical FLOP counts, dominate NPU performance. Operators must be co-designed with: (1) systolic-compatible dataflow, (2) predictable memory access, and (3) minimized DMA transfers to approach effective hardware limits.
V Discussion
Extending our findings to extreme-scale inference, we identify critical co-design considerations:
[…]
SHAVE Core Bottlenecks in Element-wise Operations While NPUs excel at matrix multiplication via systolic arrays, element-wise operations (e.g., $\operatorname{softmax}$, $\operatorname{scaling}$) execute on general-purpose SHAVE cores. At $N>1024$, these operations dominate latency in recurrent attention variants (up to 76% utilization). Kernel fusion or dedicated acceleration is essential.
DMA Management for Memory-Intensive Ops Tensor concatenation and state management consume 40-50% of cycles in Fourier attention (Table II). DMA overheads stem from frequent allocation/deallocation of large buffers. Offloading these operations to the CPU reduces latency by 32% in our tests.
[…]
VI Related Works
VI-A Long-Context Inference on Edge Platforms
Prior work has explored deploying transformer-based causal large language models (LLMs) on edge platforms [13], including hardware-specific optimizations for ARM CPUs [14] and FPGA-based execution through frameworks like llama.cpp [15, 16]. While these efforts target on-device inference, they are not designed for long-context scenarios and do not address the associated memory and compute bottlenecks that emerge in attention-heavy models. Our work addresses this gap by empirically analyzing a range of causal inference mechanisms—including standard transformers and structured state-space models (SSMs)—under long-context settings. This enables us to derive architectural insights that inform the co-design of attention mechanisms for Neural Processing Units (NPUs), supporting more efficient hardware-aware deployment strategies.
VI-B Acceleration of Sequence Models on NPUs
Several efforts have investigated transformer acceleration on NPUs [17, 18], typically through operator-level scheduling or compiler-level block partitioning. However, these approaches fall short in capturing the fine-grained resource behavior required for efficient long-context inference. Other work has focused on optimizing SSMs for NPUs [19, 20], leveraging architectural properties such as linear recurrence and memory compression. While successful within their respective domains, these strategies are not directly applicable to transformer-style causal attention. In contrast, our approach uses execution profiling and performance modeling—grounded in structured operator variants—to analyze architectural trade-offs across attention and SSM-style models. By leveraging structured state-space duality (SSD), we characterize how causal operators interact with NPU memory and compute hierarchies, enabling more informed co-design for future inference systems.
VII Conclusion
Deploying long-context AI on edge devices centers on resolving the fundamental mismatch between NPU architectures—optimized for dense, regular computations—and the memory-intensive, irregular access patterns of quadratic attention. Our analysis reveals catastrophic hardware underutilization (>>95% stalls) in standard approaches, while demonstrating that structured sub-quadratic operators (Toeplitz, Linear) transform the bottleneck into manageable bandwidth constraints. This necessitates a paradigm shift: throughput gains come not from incremental attention optimizations, but from co-designing causal operators that respect systolic dataflows and memory hierarchies. By aligning algorithmic structure with NPU execution models, we unlock the path to pervasive, private, and powerful edge AI.